OpenStudy (tocuteforyou2):

Between which two integers does √29 lie?

OpenStudy (lunacat):

http://brainly.com/question/2188223

OpenStudy (3mar):

May I help?

OpenStudy (tocuteforyou2):

yes plz

OpenStudy (holsteremission):

Notice that\[25<29<36\]Since \(\sqrt x\) is an increasing function, this means that \[\sqrt{25}<\sqrt{29}<\sqrt{36}\implies \cdots\]

OpenStudy (tocuteforyou2):

these are the options Question 6 options: A) 4 and 5 B) 13 and 14 C) 5 and 6 D) 14 and 15

OpenStudy (3mar):

Sorry for late.

OpenStudy (tocuteforyou2):

do you know the answer

OpenStudy (3mar):

me? yes

OpenStudy (tocuteforyou2):

what is it

OpenStudy (tocuteforyou2):

help me plz

OpenStudy (3mar):

It is not helpful for you to be given the answers without understanding how they are figured out!

OpenStudy (3mar):

@HolsterEmission has explained it well!

OpenStudy (tocuteforyou2):

can you show me how to solve it

OpenStudy (3mar):

Of course. With my pleasure!

OpenStudy (tocuteforyou2):

i know 25<29<36

OpenStudy (3mar):

so do you know how it is gotten!?

OpenStudy (tocuteforyou2):

is it 14and15

OpenStudy (3mar):

"Between which two integers does √29 lie" means: what are the two perfect squares between them 29 lies! The same meaning!